### v1

# Neural Force Field: Learning Generalized Physical Representation from a Few Examples

Abstract
Physical reasoning is a remarkable human ability that enables rapid learning and generalization from limited experience. Current AI models, despite extensive training, still struggle to achieve similar generalization, especially in Out-of-distribution (OOD) settings. This limitation stems from their inability to abstract core physical principles from observations. A key challenge is developing representations that can efficiently learn and generalize physical dynamics from minimal data. Here we present Neural Force Field (NFF), a modeling framework built on Neural Ordinary Differential Equation (NODE) that learns interpretable force field representations which can be efficiently integrated through an Ordinary Differential Equation (ODE) solver to predict object trajectories. Unlike existing approaches that rely on high-dimensional latent spaces, NFF captures fundamental physical concepts such as gravity, support, and collision in an interpretable manner. Experiments on two challenging physical reasoning tasks demonstrate that NFF, trained with only a few examples, achieves strong generalization to unseen scenarios. This physics-grounded representation enables efficient forward-backward planning and rapid adaptation through interactive refinement. Our work suggests that incorporating physics-inspired representations into learning systems can help bridge the gap between artificial and human physical reasoning capabilities.
Machine Learning, ICML
1 Introduction
Physical reasoning, the ability to understand and predict how objects interact in the physical world, is fundamental to both human intelligence and artificial systems (Spelke, 2022). This capability underlies crucial applications ranging from robotics to scientific discovery, making it a central challenge in AI research (Lake et al., 2017). One of the remarkable aspects of human cognitive capabilities is the ability to rapidly learn from limited examples (Kim et al., 2020; Jiang et al., 2022; Lake & Baroni, 2023; Zhang et al., 2024), especially evident in intuitive physics (Kubricht et al., 2017; Bear et al., 2021). Humans can quickly abstract core physical principles after observing limited physical phenomena, enabling them to predict complex dynamics and interact with novel environments (Spelke & Kinzler, 2007; Battaglia et al., 2013; Bonawitz et al., 2019; Xu et al., 2021).
In contrast, current AI systems face significant limitations in physical reasoning. Despite being trained on gigantic datasets, these models still struggle to achieve human-level generalization, particularly in Out-of-distribution (OOD) settings (Lake et al., 2017; Zhu et al., 2020). The core issue lies in their tendency to overfit observed trajectories rather than capturing inherent physical principles, severely limiting their ability to compose known knowledge and predict outcomes in novel contexts (Qi et al., 2021; Li et al., 2022; Wu et al., 2023). This stark contrast between human capabilities and current model limitations has motivated the search for new approaches that can learn generalizable physical representations from minimal data.
To bridge this gap, we aim to develop agents with few-shot physical learning abilities that achieve robust generalization across diverse environments. This ambitious goal presents three fundamental challenges: (i) Diverse Physical Dynamics: Physical systems exhibit intricate and nonlinear dynamics shaped by complex object properties and interactions. OOD scenarios often present drastically different dynamics from training examples, requiring sophisticated representations that explicitly capture core physical principles. (ii) Risk of Overfitting: The few-shot learning setting dramatically increases the challenge of generalization compared to large-scale training approaches. Models must carefully balance between fitting observed examples and extracting broader physical principles. (iii) Interactive Reasoning: Effective physical reasoning demands more than passive observation—agents must actively engage with their environment through experimentation and feedback, adapting their understanding based on limited examples.
To address these challenges, we introduce Neural Force Field (NFF), a physical reasoning framework building upon Neural Ordinary Differential Equation (NODE) for efficient interactive learning and reasoning. At its core, NFF employs a neural network to learn dynamic latent force fields from external interventions and object interactions. These predicted forces are then integrated through an ODE solver to compute explicit physical variables such as velocity and displacement, producing interpretable results that align with established physical principles.
Our framework offers three advantages. First, by representing physical interactions in low-dimensional force fields, NFF can rapidly learn fundamental physical concepts from just a few training examples. Second, due to the ODE-grounded dynamic graph, NFF can effectively generalize to OOD scenarios. Third, the integration of forces through an ODE solver enables fine-grained physical interactions, supporting precise modeling of collisions, gravity effects, and real-time planning.
We validate our approach on two challenging physical reasoning benchmarks: I-PHYRE (Li et al., 2023) and N-body problems (Newton, 1833). These tasks feature complex dynamics ranging from short-range forces (collision and sliding) to long-range forces (gravity), providing a comprehensive test of our framework’s capabilities. Our experiments demonstrate that NFF not only learns dynamics efficiently by abstracting physical interactions into force fields but also achieves strong generalization in both within-scenario and cross-scenario settings. Moreover, the framework’s physics-based representation enables effective forward and backward planning in goal-directed tasks, consistently outperforming existing approaches such as Interaction Network (IN) and transformer-based methods.
2 Related work
Physical reasoning  Research in physical reasoning has progressed along two main trajectories: passive observation and interactive platforms. The passive observation approach, exemplified by the Violation-of-Expectation (VoE) paradigm (Spelke et al., 1992), evaluates physical understanding by measuring agents’ ability to detect violations of intuitive physics principles (ee Baillargeon, 1987; Hespos & Baillargeon, 2001; Dai et al., 2023). While this approach has provided valuable insights into basic physical comprehension, it is limited by its inability to assess active interventions and complex reasoning.
To enable more comprehensive evaluation, interactive platforms such as PHYRE (Bakhtin et al., 2019), the virtual tools game (Allen et al., 2020), and I-PHYRE (Li et al., 2023) have emerged. These environments require agents to actively manipulate objects to achieve specific goals, testing not only prediction capabilities but also planning and reasoning skills. However, existing approaches in these platforms often struggle with generalization across diverse scenarios, particularly in few-shot settings where limited training data is available (Qi et al., 2021; Li et al., 2023).
Current physical reasoning models face two primary challenges: the need for extensive training data and limited cross-scenario transferability. While some methods achieve strong performance within specific scenarios (Allen et al., 2020), they often fail to generalize their understanding to novel situations, falling short of human-level reasoning capabilities (Kang et al., 2024). Our work addresses these limitations by introducing a framework that unifies passive observation and interactive learning through dynamic force fields. The NFF approach enables few-shot learning of physical principles while supporting active reasoning through its interpretable force-based representation, facilitating both accurate prediction and effective intervention planning across diverse physical scenarios.
Figure 1: Comparison between traditional interaction modeling and NFF. (a) Traditional methods encode physical interactions as high-dimensional latent vectors and predict future states through learned transitions between these vectors. This approach requires extensive training data and often leads to overfitting, limiting generalization to novel scenarios. (b) Our NFF framework represents interactions as learnable force fields predicted from the dynamic object graphs, which are integrated through a differentiable ODE solver to predict trajectories. NFF enables few-shot learning of various interaction types from limited interventions while maintaining strong generalization capabilities. The framework consists of four key operators: E (encoder network for processing physical scenes), D (decoder network for state reconstruction), F (force field predictor), and $\int$ (numerical integrator for computing trajectories).
Dynamic prediction  The prediction of physical dynamics, fundamental to intuitive physics, has evolved along two methodological branches: discrete and continuous-time approaches. Discrete methods typically combine recurrent architectures with GNNs (Battaglia et al., 2016; Qi et al., 2021) or transformer modules (Wu et al., 2023) for modeling object interactions, while leveraging convolutional operators for processing pixel-based information (Shi et al., 2015; Wang et al., 2022). Despite their flexibility in handling various interaction types, these methods often struggle with two key limitations: difficulty in extracting robust physical representations and susceptibility to error accumulation over extended time horizons.
Continuous-time methods address these temporal consistency issues by incorporating physical inductive biases via explicit modeling of state derivatives within dynamical systems (Chen et al., 2018; Greydanus et al., 2019; Zhong et al., 2020; Cranmer et al., 2020; Norcliffe et al., 2020). While this approach improves long-term prediction stability, current systems typically assume energy-conservative systems with simplified dynamics. More importantly, these methods face significant challenges in few-shot learning scenarios and struggle with cross-scenario generalization (Chen et al., 2020), often requiring specific physical priors and grammars to achieve meaningful transfer (Xu et al., 2021).
Our work advances continuous-time methods by introducing a more general physical representation framework via learnable force fields. This approach enables robust cross-scenario generalization while handling complex non-conservative energy systems. By combining the stability advantages of continuous-time modeling with flexible force field representations, NFF achieves both accurate long-term predictions and strong generalization capabilities without requiring explicit physical priors.
3 Method
3.1 The Neural Force Field (NFF) representation
Traditional approaches to modeling physical interactions typically rely on implicit representations through hidden vectors to describe object interactions. These methods use neural networks to learn state transition functions that map current scene features to future states. However, this purely data-driven approach cannot guarantee physically grounded representations, leading to poor generalization despite intensive training. We introduce NFF, which addresses these limitations by learning physics-grounded, generalizable representations through explicit force field modeling; see Figure 1 for a comparison.
Prediction force field  The core of NFF is built on the physical concept of a force field—a vector field that determines the force experienced by any query object based on its state $\mathbf{z}^{q}$. For a scene with $N$ objects, we model the force field $\mathbf{F}(\cdot)$ at time $t$ using a neural network operating on a relation graph $\mathcal{G}$. The graph contains object nodes $V=\{\mathbf{z}^{0}(t),\mathbf{z}^{1}(t),...,\mathbf{z}^{N-1}(t)\}$. Following neural operator methods (Lu et al., 2021) and graph neural models (Battaglia et al., 2018), we formulate the force field function as:
$\mathbf{F}(\mathbf{z}^{q}(t))=\sum_{i\in\mathcal{G}(q)}\mathbf{W}\left(f_{\theta}(\mathbf{z}^{i}(t))\cdot f_{\phi}(\mathbf{z}^{q}(t))\right)+\mathbf{b},$ (1)
where $\mathcal{G}(q)$ represents the neighbors of query $q$, $f_{\theta}$ and $f_{\phi}$ are neural networks with parameters $\theta$ and $\phi$, and $\mathbf{W}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{force}}},\mathbf{b}\in\mathbb{R}^{d_{\text{force}}}$ map hidden features to low-dimensional forces. The state vector $\mathbf{z}$ incorporates zero-order (position, angle) and first-order (velocity, angular velocity) states. This representation naturally handles various physical interactions including collision, friction, spring forces, and rotation.
Computing trajectory with ODE Rather than relying on neural decoders, NFF employs a second-order ODE integrator to compute object trajectories from the learned force field. This approach ensures physically consistent trajectories governed by fundamental principles. The dynamic change of motion for a query state follow:
$\frac{d^{2}x^{q}(t)}{dt^{2}}=\frac{dv^{q}(t)}{dt}\propto\mathbf{F}(\mathbf{z}^{q}(t)).$ (2)
While object mass theoretically affects acceleration, these mass constants are absorbed into the learned $F(\cdot)$.
Solving ODEs We solve the ODE system using numerical integration methods such as Runge-Kutta:
$\mathbf{z}^{q}(t)=\text{ODESolve}\left(\mathbf{z}^{q}(0),\mathbf{F},0,t\right),$ (3)
[…]
The NFF framework offers three key advantages over existing approaches. First, it enables few-shot learning of force fields from minimal examples—even single trajectories—through its compact, physically-grounded representation $\mathbf{F}(\mathbf{z}^{q}(t))$. Second, by representing interactions as composable force fields, NFF naturally generalizes to novel scenarios with varying object configurations, interaction types, and time horizons. As shown in Equation 1, this compositionality enables transfer of learned force patterns across physical scenarios through simple force summation. Third, the high-precision integration decoder, formulated in Equations 3 and 4, enables both fine-grained modeling of continuous-time interactions and efficient optimization for planning tasks through its invertible information flow. This precision supports various applications, from determining initial conditions for desired outcomes to designing targeted force fields for specific behaviors. The effectiveness of these three components—low-dimensional representation, ODE grounding, and precise integration—is systematically validated through ablation studies in Section 4.5.
Training NFF The framework employs autoregressive prediction, using previously predicted states to generate current states. Network optimization minimizes the Mean Squared Error (MSE) loss between predictions and ground truth. To promote learning of local, generalizable dynamics rather than global patterns, we segment long trajectories into smaller units during training. Our experiments demonstrate that this approach significantly improves convergence.
(a) I-PHYRE sim (Li et al., 2023)
(b) N-body sim (Rein & Liu, 2012)
(c) Falling
(d) Sliding
(e) Collision
(f) Spring
(g) Visualization of learned physical concepts
(h) Gravitational field ground truth
(i) Learned gravitational field
Figure 2: Visualization of learned physical dynamics and force fields. (a) I-PHYRE simulation environment demonstrating complex rigid body interactions. (b) N-body gravitational simulation showing orbital trajectories. (c) Examples of learned physical concepts demonstrated through trajectory predictions: falling under gravity, sliding with friction, collision with momentum transfer, and spring-like elastic interactions. (d-e) Comparison between ground truth and learned gravitational field distributions shows accurate force field reconstruction. Our NFF successfully captures these fundamental physical behaviors from few-shot learning.
(a) Initial configuration
(b) Ground truth trajectory
(c) Our NFF prediction
(d) SlotFormer (Wu et al., 2023)
(e) IN (Battaglia et al., 2016)
Figure 3: Trajectory predictions on unseen scenarios after few-shot learning. We compare model predictions across four physical scenarios: (i) A seesaw mechanism launching a ball over a wall, demonstrating complex contact dynamics and energy transfer, (ii) A path-building puzzle where white blocks can be eliminated during interaction, requiring the ball to roll across gaps, (iii) A dual-comet gravitational slingshot shows orbital deflection around a massive celestial body, and (iv) An eight-body gravitational system emulating solar system exhibits complex orbital dynamics around a central mass (i.e., Sun). Light trails indicate object trajectories over time. Notably, our NFF predictions closely match the ground truth behaviors across these diverse scenarios, from rigid body interactions to gravitational dynamics. Additional visualizations are provided in Appendix F.
3.2 Interactive planning
NFF extends beyond forward prediction to enable mental simulation for planning tasks. Given goal-directed scenarios, NFF can serve as a learned simulator to either search for action sequences that achieve the desired goal state through forward simulation, or optimize initial conditions and system parameters through backward computation.
Forward planning  For physical reasoning tasks requiring specific action sequences, we extend the NFF framework to incorporate action effects by including action features in $f_{\theta}$. Through forward simulation with the NFF model, we sample multiple action sequences $A=\{a_{0},...,a_{T}\}$ and optimize according to evaluation metrics $R$:
[…]
The optimal sequence $A^{\star}$ can then be executed in the physical environment. While the forward simulation may initially deviate from actual observations, NFF’s physics-based design enables efficient adaptation through new experimental data. This capability for rapid model refinement mirrors human reasoning (Allen et al., 2020) and overcomes the catastrophic forgetting challenges faced by pure neural network approaches (Dohare et al., 2024). When executed trajectories deviate from the desired goal state, the new state sequences can be directly incorporated into model optimization through the MSE loss, following the same training procedure used in the initial learning phase.
Backward planning  By inverting the ODE integrator, NFF enables the computation of initial conditions given a desired goal state. Through backward integration, the initial conditions can be determined analytically:
[…]
The invertible nature of the NFF formulation makes this backward computation particularly efficient. By performing temporal integration in reverse, we obtain initial conditions consistent with both the goal state and the learned physical dynamics encoded in the NFF model.
(a) Support interaction
(b) Platform orientation
(c) Impact dynamics
(d) Spring mechanics
Figure 4: Force response analysis under controlled state variations. Each plot shows the predicted force components (Fx in blue, Fy in orange) from our trained NFF model, measured in normalized units against different state parameters. The subplots systematically analyze: (a) support forces as a function of ball height, showing characteristic contact response when the ball meets the platform, (b) force decomposition as the platform rotates from horizontal ($0^{\degree}$) to vertical ($90^{\degree}$), (c) impact forces scaling with the ball’s downward velocity, and (d) spring forces varying with horizontal displacement from the equilibrium position. Each plot varies one parameter while keeping all other scene variables constant, demonstrating NFF’s learned physical principles, including contact mechanics, geometric reasoning, impact dynamics, and harmonic motion.
4 Experiments
4.1 Environments and datasets
We evaluate our approach on two physical reasoning tasks: I-PHYRE and N-body dynamics. Each task is evaluated across three distinct settings: within-scenario prediction, cross-scenario prediction, and planning. Detailed environment and dataset configurations are provided in Appendix A.
I-PHYRE   I-PHYRE (Li et al., 2023) presents a suite of complex physical reasoning puzzles requiring multi-step interventions. The environment incorporates diverse physical interactions including gravity, collision, friction, rotation, spring dynamics, and pendulum motion. It challenges AI agents to solve puzzles with minimal environmental interactions while generalizing to unseen scenarios. For within-scenario prediction, we evaluate on 10 training games that share similar scenarios but require different solutions. The cross-scenario prediction setting extends to 30 novel games featuring noise, compositional elements, and multi-ball scenarios, with varying object properties such as block lengths, ball sizes, and object positions. In the planning setting, models must generate optimal action sequences to successfully complete each game.
N-body  The N-body problem (Newton, 1833) tests trajectory prediction for small comets orbiting a massive central planet. Using REBOUND simulator (Rein & Liu, 2012), we generate dynamics data by randomly sampling orbital parameters including radii, angles, and masses. This task evaluates the model’s ability to infer gravitational laws from limited observations. The within-scenario prediction setting introduces novel initial conditions and masses, focusing on systems with 1-2 orbiting bodies tracked over 50 timesteps. Cross-scenario prediction significantly increases complexity by introducing systems with 8 orbiting bodies tracked over 100 timesteps. The planning setting challenges models to optimize initial conditions that will evolve to specified target states after 50 timesteps.
4.2 Learning force fields from a few examples
We qualitatively evaluate NFF’s ability to learn force fields from limited training data; training implementation details are provided in Appendix B. For I-PHYRE, we train NFF on just 10 basic games with 10 trajectory samples each. From these 100 trajectories, the model successfully learns fundamental physical concepts including gravity, support, sliding, collision, friction, and spring dynamics, representing them through unified force fields (Figure 2(g)). To verify that NFF learns generalizable physical principles, we examine force responses to varied ball-block interactions under controlled conditions. Figure 4 shows our systematic evaluation where we independently vary individual scene parameters such as position, velocity, and angle while holding others constant, demonstrating NFF’s accurate force response modeling.
For the N-body system, we train NFF using 200 randomly sampled trajectories from 2-body and 3-body dynamics. As shown in Figure 2(i), the model successfully learns to capture the inverse gravitational field, correctly modeling the distance-dependent centripetal forces governing the mutual attraction between bodies.
(a) I-PHYRE
(b) N-body
Figure 5: NFF outperforms other baselines in prediction tasks. Performance comparison between IN (Battaglia et al., 2016), SlotFormer (Wu et al., 2023), and our NFF on (a) I-PHYRE and (b) N-body tasks. Box plots display median, quartiles, and outliers, with lower values $\downarrow$ better for error metrics (RMSE, FPE, PCE) and higher values $\uparrow$ better for correlation (R). Our NFF consistently achieves lower errors and higher correlations across both simulation environments, particularly in challenging cross-scenario generalization.
4.3 Prediction on unseen scenarios
We evaluate the learned force fields’ predictive accuracy in both within-scenario and cross-scenario settings. For I-PHYRE, we test against 20 ground truth trajectories per game, while for N-body systems, we evaluate using 200 novel initial conditions. We compare our results against established interaction modeling methods: IN (Battaglia et al., 2016) and SlotFormer (Wu et al., 2023).
Evaluation metrics  We employ multiple complementary metrics to assess prediction quality. Beyond standard Root Mean Squared Error (RMSE), we introduce Final Position Error (FPE) (final position error) and Position Change Error (PCE) (position change error) for detailed performance analysis. FPE quantifies terminal position accuracy, while PCE evaluates the model’s ability to capture motion dynamics. Additionally, Pearson Correlation Coefficient (R) measures trajectory shape alignment independent of speed variations. This comprehensive metric suite enables thorough evaluation across temporal and spatial dimensions. Detailed definitions are provided in Section B.2.
Results  As shown in Figure 3, NFF generates physically plausible trajectories that closely match ground truth behavior, even in previously unseen scenarios. Quantitative comparisons in Figure 5 demonstrate NFF’s superior performance across all metrics for both I-PHYRE and N-body tasks, with particularly strong results in cross-scenario generalization. While SlotFormer exhibits overfitting tendencies that limit its cross-scenario performance (analyzed further in Appendix C), NFF’s ability to learn generalizable physical principles from limited training data enables robust prediction across diverse scenarios.
4.4 Planning on unseen scenarios
The trained NFF model can generate plans for novel tasks after learning from limited demonstrations. Unlike prediction tasks that evaluate trajectory accuracy, planning tasks require generating action sequences to achieve specific goals.
I-PHYRE planning  We implement a 5-round interactive learning protocol. NFF acts as a mental simulator to evaluate 500 randomly sampled action candidates, selecting the optimal sequence for physical execution. After each execution, the model updates its parameters based on observed outcomes, refining its physics understanding. This updated model then guides subsequent action proposals.
Figure 6: Interactive planning performance on I-PHYRE. Comparison of cumulative success probability over 5 planning rounds between different approaches: human performance random sampling, our NFF, IN, and SlotFormer. Solid and dashed lines indicate variants with and without the refinement mechanism (Allen et al., 2020), respectively. Error bars show SEM across trials. Our NFF with refinement (yellow) demonstrates continuous improvement across rounds, achieving performance comparable to human players (data from Li et al. (2023)), while baseline models show limited performance even with refinement.
(a) Euler step size
(b) Step size $25e-$ 2
(c) Step size $35e-$ 3
(d) Latent dimension
(e) Ground-truth
(f) Dim = 3
(g) Dim = 16
(h) Dim = 256
(i) Latent space
Figure 7: Ablation studies on key model components in N-body systems. Results averaged across test cases and five random seeds. (a) Model error vs. integration step size ($31e-$ 3 to $25e-$ 2), showing improved accuracy with finer Euler integration. (b-c) Performance comparison with and without ODE integration for within- and cross-scenario generalization, evaluated at different step sizes. (d) Impact of latent space dimensionality (3, 16, 256) on generalization, demonstrating optimal performance when matching the inherent 3D nature of gravitational forces. (e) UMAP visualization of learned force field representations, revealing that lower dimensionality better captures the underlying gravitational field manifold compared to higher dimensions. Blue/orange indicates relative magnitude of projected interactions.
We quantitatively evaluate planning performance by calculating success probability $p_{i}$ for each game as the success rate over 20 trials in round $i$, with NFF updating after failures; see detailed results in Appendix E. Figure 6 compares NFF against random sampling, human performance (from Li et al. (2023)), IN, and SlotFormer using cumulative success probability: $1-\prod_{i=1}^{n}(1-p_{i})$ for each trial $n$. NFF outperforms baselines and approaches human-level performance after refinement (Allen et al., 2020), demonstrating effective few-shot learning. The poor performance of IN and SlotFormer, falling below random sampling, indicates how inaccurate dynamics modeling compromises planning.
N-body planning  We focus on determining initial conditions that achieve desired final configurations under celestial dynamics. NFF’s inverse simulation capability enables direct computation of initial conditions through reverse time evolution, while IN and SlotFormer rely on iterative gradient-based refinement. As shown in Table 1, NFF achieves superior planning performance in both within-scenario and cross-scenario settings.
Table 1: N-body system initial condition reconstruction from target configurations. MSE $\downarrow$ comparison between IN, SlotFormer, and our NFF. Results show average and SEM across trials for both within- and cross-scenario evaluations. Lower values indicate better reconstruction accuracy.
Scenario IN SlotFormer NFF
Within $\downarrow$ 0.651 $\pm$ 0.021 0.837 $\pm$ 0.029 0.067 $\pm$ 0.010
Cross $\downarrow$ 4.654 $\pm$ 0.193 4.018 $\pm$ 0.133 0.140 $\pm$ 0.011
4.5 Ablation study
We investigate three key factors affecting model performance: integration precision, ODE grounding, and latent dimensionality. For integration precision, we evaluate NFF on the N-body task using Euler integration at four precision levels: $31e-$ 3, $35e-$ 3, $21e-$ 2, and $25e-$ 2. Figure 7(a) demonstrates that higher integration precision consistently yields better performance. We also explore more integration methods in Appendix D. To study ODE grounding’s impact on generalization, we compare two configurations: our NFF using predicted force fields with ODE integration, and IN using learned state transitions at matching step sizes. Figures 7(b) and 7(c) shows that ODE integration enhances generalization performance. However, at coarser step sizes ($25e-$ 2), NFF slightly underperforms IN, indicating that ODE-based approaches require fine integration precision to realize their advantages. For latent dimensionality analysis, we vary the latent interaction space dimension in NFF across 3, 16, and 256 dimensions. Results in Figure 7(d) reveal that lower-dimensional representations improve generalization capability. The visualization in Figure 7(i) confirms this finding, showing that reduced dimensionality leads to latent representations more closely matching ground truth. These ablation studies demonstrate that NFF achieves optimal few-shot learning of physical dynamics when combining high integration precision, low-dimensional latent space, and ODE-based integration.
5 Discussion
Modeling diverse forces  While NFF has demonstrated effectiveness in modeling fundamental physical forces like gravity, support, and collision, its principles extend naturally to broader domains. The framework can potentially model social forces (Shu et al., 2015; Xie et al., 2017; Wei et al., 2017, 2018; Shu et al., 2021) and biological forces governing migration patterns and human motion (Netanyahu et al., 2024). Limitations  Like existing approaches (Battaglia et al., 2016), NFF operates on symbolic states, setting aside perception challenges. The integration of raw sensory input with physics learning remains an open challenge, requiring investigation into various representation frameworks.
6 Conclusion
We present NFF, a force field-based representation framework for modeling physical interactions that exhibits human-like few-shot learning, generalization, and reasoning capabilities. Our experiments on two challenging physical reasoning datasets demonstrate NFF’s ability to learn diverse physical concepts and rules from limited observations. This initial exploration of force field opens new possibilities for developing physical world models through representation learning, potentially bridging the gap between symbolic physics understanding and learning-based approaches.
Impact statements
This work contributes to advancing physics-informed machine learning with potential applications in robotics, simulation, and scientific discovery. The proposed NFF framework could enable more sample-efficient and interpretable physical modeling, reducing computational resources needed for training. While primarily focused on fundamental physics, the methodology could extend to social and biological systems modeling. We acknowledge potential risks if misused for simulating harmful physical scenarios. However, the research focuses on basic physical principles and publicly available phenomena, prioritizing transparent and reproducible science. The computational requirements for training NFF are moderate, making it accessible to most research institutions while maintaining a reasonable carbon footprint.
Acknowledgements
The authors would like to thank Miss Chen Zhen (BIGAI) for making the nice figures, Prof. Yizhou Wang (Peking University) for the helpful discussion, and Minyang Yu (Peking University) for working on the GPT’s results. S. Li, R. Shen, and Y. Zhu are supported in part by the National Natural Science Foundation of China (62376031), the Beijing Nova Program, the State Key Lab of General AI at Peking University, the PKU-BingJi Joint Laboratory for Artificial Intelligence, and the National Comprehensive Experimental Base for Governance of Intelligent Society, Wuhan East Lake High-Tech Development Zone.
Appendix A Environment and dataset configurations
A.1 I-PHYRE
We adopt the original settings from the I-PHYRE work (Li et al., 2023). The training dataset consists of 10 basic games: support, hinder, direction, hole, fill, seesaw, angle, impulse, pendulum, and spring. For each game, we randomly generate 5 successful and 5 failed action sequences. The within-scenario setting includes an additional 10 successful and 10 failed action sequences for each of the 10 basic games. The cross-scenario setting contains 10 successful and 10 failed action sequences from 30 unseen games, which include 10 noisy games, 10 compositional games, and 10 multi-ball games. All object masses are set to be equal, and small friction and elasticity coefficients are applied. For each game, we record the center positions, lengths, angles, radii, horizontal velocities, vertical velocities, and rotational velocities of each object, as well as the spring pair indices. Each sequence contains up to 12 objects and spans 150 timesteps.
A.2 N-body
We employed the REBOUND engine to simulate N-body dynamics and considered two types of N-body problems for 3D trajectories: the planetary n-body problem and the cometary n-body problem. In the planetary n-body problem, the comets are initialized with Keplerian velocities, while in the cometary n-body problem, the planets are initialized with escape velocities. Initial positions are sampled from the spherical coordinate system and then converted to Cartesian coordinates. The radii are sampled within the range of 1 to 3, azimuthal angles are sampled from 0 to $2\pi$, and polar angles are sampled from $-\pi/6$ to $\pi/6$. Masses are sampled from 0.05 to 0.1, while the central massive body is assigned a mass of 5. All sampling is performed using Latin Hypercube Sampling to ensure comprehensive space coverage, even with sparse samples. The training dataset includes 50 two-body planetary trajectories, 50 three-body planetary trajectories, 50 two-body cometary trajectories, and 50 three-body cometary trajectories, spanning across 50 timesteps. The within-scenario dataset consists of the same four trajectory types as the training set but with different initial conditions. The cross-scenario dataset contains 100 nine-body planetary trajectories with radii ranging from 1 to 11 and 100 nine-body cometary trajectories with radii ranging from 1 to 5. The trajectory length in the cross-scenario data is extended to 100 steps. We record the masses, 3D Cartesian coordinates, and velocities for training.
Appendix B Training details
B.1 Hyperparameters
In I-PHYRE, we train the models using a single NVIDIA A100 Tensor Core GPU. The force field predictor in NFF is based on a DeepONet architecture, consisting of trunk and branch networks, each a 3-layer MLP with a hidden size of 256. The ODE solver uses the Euler method with a step size of 0.005. The 150-step trajectories are divided into 6-step segments, with 6 trajectories per batch. The learning rate starts at 5e-4 and gradually decays to 1e-5 following a cosine annealing schedule. Training occurs over 3000 epochs with a weight decay of 1e-5 for regularization.
The interaction predictor in IN is a 3-layer MLP with a hidden size of 256, while the decoder is also an MLP of the same size. SlotFormer uses 3 transformer encoder layers, each with 4 attention heads and a feedforward network of size 256. Both IN and SlotFormer are trained with a batch size of 50 trajectories, employing the same segmentation technique. All other hyperparameters are kept consistent with those used in NFF.
In N-body, we train the models using a single NVIDIA GeForce RTX 3090 GPU. For all models, the 50-step trajectories are divided into 5-step segments, with 50 trajectories per batch. The learning rate starts at 5e-4 and gradually decays to 1e-7 following a cosine annealing schedule. Training occurs over 5000 epochs with a weight decay of 1e-5 for regularization. All the other hyperparameters are the same as those used in I-PHYRE.
B.2 Evaluation metrics
In this study, the chosen evaluation metrics include, RMSE, FPE, PCE, and R. Each of these metrics provides valuable insights into different characteristics of the model’s predictions, allowing for a comprehensive evaluation.
Root Mean Squared Error (RMSE)
The RMSE, the square root of MSE, shares the same characteristics as MSE in terms of penalizing larger errors. However, it provides the error in the same units as the original data, allowing for a more intuitive understanding of how far off the model’s predictions are, on average, in the context of physical trajectories:
RMSE$=\sqrt{\frac{1}{n}\sum_{t=1}^{n}(\hat{z}_{t}-z_{t})^{2}}.$ (A1)
Final Position Error (FPE)
The FPE specifically measures the discrepancy between the predicted and actual final position of the object at the end of the trajectory. This metric is crucial for goal-driven physical reasoning tasks where the objects are expected to finally move into a specific area. FPE helps ensure that the model is not only capturing the intermediate trajectory but also predicting the final destination with high accuracy:
FPE$=\left|\hat{z}_{\text{final}}-z_{\text{final}}\right|.$ (A2)
Position Change Error (PCE)
The PCE quantifies the error in the predicted change of position over time, which can be interpreted as a measure of the model’s accuracy in tracking the object’s velocity during its motion:
PCE$=\left|\Delta\hat{z}_{t}-\Delta z_{t}\right|.$ (A3)
Pearson Correlation Coefficient (R)
The R measures the linear relationship between the predicted and actual trajectories. R is useful for assessing how well the model captures the overall trend or pattern of the trajectory, even when the absolute errors might vary. A high correlation suggests that the model is effectively tracking the overall movement, while a low correlation might indicate that the model fails to capture the underlying trajectory pattern:
R$=\frac{\sum_{t=1}^{n}(\hat{z}_{t}-\bar{\hat{z}})(z_{t}-\bar{z})}{\sqrt{\sum_{t=1}^{n}(\hat{z}_{t}-\bar{\hat{z}})^{2}\sum_{t=1}^{n}(z_{t}-\bar{z})^{2}}}.$ (A4)
Appendix C Overfitting
Figure A1: SlotFormer overfitting on physical data. While the performance improves in the within-scenario setting as training progresses, the cross-scenario setting experiences overfitting, resulting in increasing RMSE during training.
We demonstrate that methods without interaction modeling are prone to overfitting on in-distribution physical data. While they perform well within the same scenario, they face significant challenges in cross-scenario settings. To illustrate this, we train a SlotFormer on N-body training data and validate it on a separate 4-body cross-scenario dataset. As shown in Figure A1, the increase in overfitting leads to a decline in generalization performance.
Appendix D Ablation studies on integration methods
In Table A1, we present additional ablation studies comparing various integration methods. Specifically, we evaluate the performance of different integration orders on the N-body task, including Euler, Midpoint, Heun3, RK4, and adaptive methods. Computational complexity is measured by the average number of integration steps. The results indicate that Euler integration provides the best cross-scenario generalization, while Heun3 excels in within-scenario generalization. Despite this, Euler integration has lower computational complexity than the higher-order methods. These findings suggest that Euler integration is sufficient for our physical reasoning tasks.
Table A1: Results of different kinds of integration methods. The computational complexity of the model is proportional to the number of average integration steps (relative to Euler $35e-$ 3). Among non-adaptive integration methods, the Euler method achieves the best generalization performance with relatively low computational complexity, whereas adaptive methods at different tolerance levels do not exhibit consistent performance improvements.
Method Average steps (relative) Within Cross
Adaptive methods (tolerance)
Adaptive ($31e-$ 3) $0.20$ $0.0725\pm 0.0084$ $0.1537\pm 0.0069$
Adaptive ($41e-$ 4) $0.28$ $0.0582\pm 0.0035$ $0.1797\pm 0.0125$
Adaptive ($51e-$ 5) $0.36$ $0.0783\pm 0.0056$ $0.1539\pm 0.0056$
Adaptive ($61e-$ 6) $0.56$ $0.0807\pm 0.0053$ $0.1668\pm 0.0067$
Adaptive ($71e-$ 7) $1.07$ $0.0811\pm 0.0053$ $0.1928\pm 0.0077$
Non-adaptive methods (step size)
Euler ($35e-$ 3) $1.00$ $0.0785\pm 0.0069$ $0.1522\pm 0.0066$
Midpoint ($35e-$ 3) $2.00$ $0.0671\pm 0.0053$ $0.1762\pm 0.0071$
Heun3 ($35e-$ 3) $3.00$ $0.0606\pm 0.0041$ $0.1550\pm 0.0067$
RK4 ($35e-$ 3) $4.00$ $0.0682\pm 0.0048$ $0.1626\pm 0.0072$
Appendix E Detailed planning results
We present the detailed planning results of different models across trials in Table A2.
Besides those dynamic prediction models, we also benchmark the planning performance of the large language model GPT-4o with the refinement mechanism. The core idea is utilizing In-Context Learning (ICL) abilities and high-level reasoning abilities of large language models to find correct solutions.
The prompt in Round 1 consists of 3 parts: (1) I-PHYRE description and an elaborately explained example. (2) Other training examples. (3) Description of the testing game setting. In Part 2, we provide GPT-4o with 1 successful solution and 1 wrong solution together with resulting trajectories of objects for each of the 10 basic games. In the refinement stage (i.e., Round 2 and Round 3), we run the action sequence proposed by GPT-4o on the simulator, Then we offer GPT-4o the resulting trajectory and ask it to modify the previous solution or design a new one.
We test the performance of GPT-4o for both within-scenario planning and cross-scenario planning. Detailed results are presented in Table A3. Within-scenario games contains the same 10 games used in the prompt in Round 1. So the language model can memorize corresponding solutions provided in the prompt to achieve good performance. Cross-scenario games contains the other 30 game settings, on which the performance of GPT-4o is better than other baselines (Random, IN, SlotFormer) in Table A2. The improvement of performance through refinement is witnessed in both within-scenario and cross-scenario games.
The prompt used in Round 1 is shown below.
Table A2: Planning results in I-PHYRE. Average probability of succeeding cross-scenario games after $n$ trails from IN, SlotFormer, NFF, human, w/o refining. The gray line indicates the best results among all the AI methods.
Method Refine Round 1 Round 2 Round 3 Round 4 Round 5
Random - $0.24\pm 0.04$ $0.37\pm 0.05$ $0.45\pm 0.06$ $0.51\pm 0.06$ $0.56\pm 0.06$
IN $\times$ $0.25\pm 0.06$ $0.33\pm 0.07$ $0.39\pm 0.07$ $0.42\pm 0.08$ $0.45\pm 0.08$
IN $\checkmark$ $0.25\pm 0.06$ $0.39\pm 0.07$ $0.48\pm 0.07$ $0.57\pm 0.07$ $0.63\pm 0.06$
SlotFormer $\times$ $0.30\pm 0.07$ $0.36\pm 0.08$ $0.39\pm 0.08$ $0.41\pm 0.08$ $0.43\pm 0.08$
SlotFormer $\checkmark$ $0.30\pm 0.07$ $0.42\pm 0.08$ $0.49\pm 0.07$ $0.55\pm 0.07$ $0.60\pm 0.07$
NFF $\times$ $0.51\pm 0.08$ $0.55\pm 0.08$ $0.58\pm 0.08$ $0.60\pm 0.08$ $0.62\pm 0.08$
NFF $\checkmark$ $0.51\pm 0.08$ $0.62\pm 0.08$ $0.69\pm 0.07$ $0.74\pm 0.06$ $0.77\pm 0.06$
Human $\times$ $0.52\pm 0.06$ $0.66\pm 0.06$ $0.73\pm 0.06$ $0.77\pm 0.06$ $0.80\pm 0.05$
Human $\checkmark$ $0.52\pm 0.06$ $0.70\pm 0.06$ $0.79\pm 0.05$ $0.83\pm 0.04$ $0.86\pm 0.04$
Table A3: GPT-4o’s planning results in I-PHYRE. The GPT-4o refines itself after each round of play and shows increasingly better performance.
Scenario Round 1 Round 2 Round 3
Within $0.74\pm 0.08$ $0.78\pm 0.07$ $0.82\pm 0.07$
Cross $0.35\pm 0.07$ $0.45\pm 0.08$ $0.51\pm 0.04$
/* PART 1 */
This is a 2D physics simulation environment. There will be one or few red balls, and may appear black blocks, grey blocks and blue blocks. Lengths of blocks are different.
All these objects are rigid bodies. When we talk about physical laws, we consider gravity (there is gravity in this environment), friction (but the friction coefficients are small), collision between rigid bodies, rotation of balls and blocks.
However, grey blocks and black blocks are fixed and stay still to their initial positions. They do not apply any laws of physics (even collision wont affect them). The red ball and blue blocks are controlled by physical laws.
Initially, all objects are at rest (their velocity and angular velocity are all 0).
There may exist 2 types of constraints between two objects: spring or joint. "Spring" constraint means there is a spring linking 2 objects. Springs in all the games have the same rest length, stiffness and damping rate (You can infer these properties from trajectories provided below that contain springs). "Joint" constraint means there is a rod supporting 2 objects so their distance is fixed.
We emphasize that grey and black blocks are not affected by any constraints. They are fixed to their initial positions.
Now, we are going to play a game. The goal is to make all the red balls fall down to the bottom of the canvas.
Lets first define a coordinate on the canvas. Let the origin be the top left corner. The x-axis is horizontal, from left to right. The y-axis is vertical, from top to bottom. The canvas size is (600, 600), so the coordinate of the lower-right corner is (600, 600).
The only operation the player can do is to eliminate any grey blocks at any time they want (however, we limit the whole game in 15 seconds. If after 15s the red ball still doesnt reach the bottom, the player loses the game). The player cannot eliminate other objects, including the red ball, blue or black blocks. Note that if the player eliminates a grey block which is connected to some other object with a spring or joint, then the spring/joint between them will disappear.
Now, I will show you a game named "fill". The name is actually the key to solve this problem. In this game, a blue block is placed above a grey block. You need to eliminate the grey block so that the blue block falls down to fill the pit on the "ground" (the ground and pit are built by a few black blocks). Otherwise, the red ball will roll into and get trapped in the pit.
Below is the initial setting of the game. I will describe the meaning of each parameter.
{’block’: [[[100.0, 420.0], [400.0, 420.0]], [[100.0, 400.0], [250.0, 400.0]], [[350.0, 400.0], [400.0, 400.0]], [[270.0, 200.0], [330.0, 200.0]], [[270.0, 180.0], [330.0, 180.0]], [[100.0, 150.0], [180.0, 200.0]], [[150.0, 100.0], [150.0, 150.0]]], ball’: [[120.0, 120.0, 20.0]], eli’: [0, 0, 0, 1, 0, 1, 1, 0], dynamic’: [0, 0, 0, 0, 1, 0, 0, 1]}
First comes the block items. Each item is in the form of ‘[[left_x, left_y], [right_x, right_y]]‘, namely, the coordinates of the left and right corner of the block. Note that each block is generated by moving a circle (radius = 10) centered from the left corner to the right corner.
Next comes the ball items. Now we only have 1 ball in this game. This item looks like ‘[x, y, r]‘, namely, the coordinates of the center and the radius of the ball.
Then comes the eli item specifying which of these objects are eliminable. In this game, eli’: [0, 0, 0, 1, 0, 1, 1, 0]. This is an 8-dimensional vector. The first 7 elements describes the 7 blocks (there are 7 elements in the block item), and the last element describes the ball. 0 means the object is not eliminable, and 1 means the object is eliminable.
Finally comes the dynamic item, which specifies whether the objects are dynamic. In this game, dynamic’: [0, 0, 0, 0, 1, 0, 0, 1]. The first 7 elements are the 7 blocks, and the last element is the ball. 0 means the object is static (not apply physical laws), and 1 means the object is dynamic (apply physical laws).
You can infer from eli and dynamic that the 4th, 6th, and 7th blocks are grey, the 5th block is blue. All other blocks are black.
In other game settings, there may exist spring or joint items. They look like spring’: [[6, 7]], which means there is only 1 spring connecting the 6th and 7th object (object can be blocks or balls). joint’: [[6, 7]] means there is only 1 joint connecting the 6th and 7th object.
Here, we also provide an image of the game.
<image here>
Now you are provided with a successful action sequence that can solve the game. The action sequence is as follows:
[[150.0, 125.0, 0.3], [300.0, 200.0, 1.4000000000000004]]
Action sequence is a list of eliminations [pos_x, pos_y, t], where pos_x and pos_y must be the center of some block (formally, the center is the average of the coordinates of 2 corners: pos_x = 1 / 2 * (left_x + right_x), pos_y = 1 / 2 * (left_y + right_y)), and t is the time when the block is eliminated (t should be in 0.1 ~ 15.0). In this example, there are 2 grey blocks to eliminate. [150, 125, 0.3] means eliminate the vertical block (next to the red ball) at t = 0.3s; [300.0, 200.0, 1.4] means eliminate the horizontal block (which supports the blue block from falling down) at t = 1.4s.
In this setting, all positions that can be eliminated are [[300.0, 200.0], [140.0, 175.0], [150.0, 125.0]].
Now you are provided with the trajectory of the red ball and the dynamic (blue) block. Other blocks are either static (the same as in the initial setting) or eliminated at some time. In this case, there is only one blue block. If there are multiple blue blocks, their position at some timestamp will be given by the same order as in dyn item. If there are no blue balls, the dynamic item will be an empty list.
Below is the trajectory.
t = 0.000s: {’ball’: [[120.0, 120.0]], dynamic_block’: [[270.0, 180.0, 330.0, 180.0]]}
t = 0.167s: {’ball’: [[120.0, 121.25]], dynamic_block’: [[270.0, 180.03, 330.0, 180.03]]}
t = 0.333s: {’ball’: [[120.0, 125.28]], dynamic_block’: [[270.0, 180.03, 330.0, 180.03]]}
/* PART 2 */
Now we provide you with more examples from different game settings. We will provide 10 game settings. For each game setting, we will provide you with the initial setting (data + image), and 2 action sequence & trajectory (including 1 successful and 1 failed ones).
Game 1: name:angle
Initial setting: {’block’: [[[100.0, 400.0], [400.0, 400.0]], [[200.0, 350.0], [210.0, 350.0]], [[200.0, 300.0], [210.0, 300.0]], [[480.0, 400.0], [550.0, 400.0]], [[390.0, 350.0], [400.0, 350.0]], [[390.0, 300.0], [400.0, 300.0]], [[100.0, 380.0], [100.0, 360.0]], [[550.0, 380.0], [550.0, 360.0]], [[200.0, 280.0], [400.0, 280.0]]], ball’: [[300.0, 250.0, 20.0]], eli’: [0, 1, 1, 0, 1, 1, 0, 0, 0, 0], dynamic’: [0, 0, 0, 0, 0, 0, 0, 0, 1, 1]}
Eliminable positions: [[205.0, 350.0], [205.0, 300.0], [395.0, 350.0], [395.0, 300.0]]
Game setting image:
<image here>
Successful Case 1:
Action sequence: [[395.0, 300.0, 1.4000000000000004], [205.0, 350.0, 2.900000000000001], [205.0, 300.0, 3.9000000000000012], [395.0, 350.0, 5.900000000000002]]
This action sequence leads to success.
Trajectory:
t = 0.000s: {’ball’: [[300.0, 250.0]], dynamic_block’: [[200.0, 280.0, 400.0, 280.0]]}
t = 0.167s: {’ball’: [[300.0, 250.32]], dynamic_block’: [[200.02, 280.3, 400.02, 280.3]]}
Failed Case 1:
Actions: [[395.0, 350.0, 3.000000000000001], [205.0, 300.0, 4.800000000000002], [205.0, 350.0, 6.400000000000002], [395.0, 300.0, 6.700000000000002]]
This action sequence leads to failure.
Trajectory:
t = 0.000s: {’ball’: [[300.0, 250.0]], dynamic_block’: [[200.0, 280.0, 400.0, 280.0]]}
t = 0.167s: {’ball’: [[300.0, 250.32]], dynamic_block’: [[200.02, 280.3, 400.02, 280.3]]}
... <more examples>
/* PART 3 */
Your goal is to provide a successful action sequence (that makes the red ball fall to the bottom of the canvas) under a new game setting. Below is the new game setting.
Game name: "angle".
Game setting:
Eliminable positions: ...
Game setting image:
<image here>
Your task is to provide a successful action sequence in the same format as examples given previously. You should make some analysis or explanations.
Now please provide one successful action sequence. At the end of your response, please give the entire action sequence in the format of
Action sequnce: [[pos_1_x, pos_1_y, t_1], [pos_2_x, pos_2_y, t_2], ...]
so that we can extract and test your proposed action sequence easily.
The prompt used in Round 1.
Appendix F More visualization
We present additional visualization of learned forces along the trajectories in I-PHYRE in Figures A2 and A3. We also present more prediction results on N-body in Figure A4.
Figure A2: The inverted forces from NFF and its predicted trajectories of dynamic objects in I-PHYRE within scenarios. Ground-truth trajectories are represented by black dots, while predicted trajectories are shown with colorful dots. Horizontal and vertical forces are depicted as arrows, while rotational forces are indicated by red or blue dots, signifying negative and positive forces, respectively. The color of the forces indicates which object is exerting the force. Only the initial states are shown for static objects. Some objects will be eliminated during the process. Best seen in videos.
Figure A3: The inverted forces from NFF and its predicted trajectories of dynamic objects in I-PHYRE cross scenarios. Ground-truth trajectories are represented by black dots, while predicted trajectories are shown with colorful dots. Horizontal and vertical forces are depicted as arrows, while rotational forces are indicated by red or blue dots, signifying negative and positive forces, respectively. The color of the forces indicates which object is exerting the force. Only the initial states are shown for static objects. Some objects will be eliminated during the process. Best seen in videos.
(a) Ground truth trajectory
(b) Our NFF prediction
(c) SlotFormer (Wu et al., 2023)
(d) IN (Battaglia et al., 2016)
Figure A4: Additional prediction results on N-body.